# Counting Stacked Objects

### v2

@@ Abstract
Visual object counting is a fundamental computer vision task underpinning numerous real-world applications, from cell counting in biomedicine to traffic and wildlife monitoring. However, existing methods struggle to handle the challenge of stacked 3D objects in which most objects are hidden by those above them. To address this important yet underexplored problem, we propose a novel 3D counting approach that decomposes the task into two complementary subproblems - estimating the 3D geometry of the object stack and the occupancy ratio from multi-view images. By combining geometric reconstruction and deep learning-based depth analysis, our method can accurately count identical objects within containers, even when they are irregularly stacked. We validate our 3D Counting pipeline on diverse real-world and large-scale synthetic datasets, which we will release publicly to facilitate further research.
@@ 1 Introduction
At the heart of our proposed solution is a key insight: The number of objects per unit of volume, which we will refer to as the occupancy ratio, can be accurately inferred from a depth map computed by a monocular depth estimator from a view in which enough objects of interest are clearly visible. In most cases, such a view is one where the container is seen roughly from above, without having to be strictly vertical. To exploit it, we break down the problem into two complementary tasks: estimating the 3D geometry of the object stack and estimating the occupancy ratio within this volume, as depicted by Fig. 2. This decomposition enables us to solve the 3D counting problem through a combination of geometric reconstruction for volume estimation and deep learning-based depth analysis for occupancy prediction, both of which can be solved efficiently.
@@ 9 Implementation details
We use the nerfstudio library [25] for 3D reconstruction, specifically the splatfacto method built on top of the gsplat library [32]. A combination of trimesh and the alphashape python library to compute alpha concave hulls and measure their volume. We thank the contributors of all the aforementioned libraries.

### v3

@@ Abstract
Visual object counting is a fundamental computer vision task underpinning numerous real-world applications, from cell counting in biomedicine to traffic and wildlife monitoring. However, existing methods struggle to handle the challenge of stacked 3D objects in which most objects are hidden by those above them. To address this important yet underexplored problem, we propose a novel 3D counting approach that decomposes the task into two complementary subproblems - estimating the 3D geometry of the object stack and the occupancy ratio from multi-view images. By combining geometric reconstruction and deep learning-based depth analysis, our method can accurately count identical objects within containers, even when they are irregularly stacked. We validate our 3D Counting pipeline on large-scale synthetic and diverse real-world datasets with manually verified total counts. Our datasets and code and can be found at https://corentindumery.github.io/projects/stacks.html
@@ 1 Introduction
At the heart of our proposed solution is a key insight: the fraction of space occupied by objects, which we will refer to as the occupancy ratio, can be accurately inferred from a depth map computed by a monocular depth estimator from a view in which enough objects of interest are clearly visible. In most cases, such a view is one where the container is seen roughly from above, without having to be strictly vertical. To exploit it, we break down the problem into two complementary tasks: estimating the 3D geometry of the object stack and estimating the occupancy ratio within this volume, as depicted by Fig. 2. This decomposition enables us to solve the 3D counting problem through a combination of geometric reconstruction for volume estimation and deep learning-based depth analysis for occupancy prediction, both of which can be solved efficiently.
@@ 5 Conclusion
6 Acknowledgements.
We would like to sincerely thank all the participants who spent time to play our counting game and provided us with valuable data on human performance at this task. In particular, we would like to congratulate Martin Engilberge, who gave the most accurate estimates and has been awarded a box of chocolates. We would also like to thank Nicolas Talabot, Víctor Batlle and Adriano D’Alessandro for insightful discussions.
@@ 7 Dataset Details
8 Additional comparison distinguishing visible and invisible objects.
Our method is, to the best of our knowledge, proposing the first solution for this task. The methods used as comparison in 4 are initially designed to count only visible objects, and perform poorly on our dataset. In this section, we distinguish between visible and invisible objects in stacks and evaluate baseline methods against both counts. To perform this evaluation, we manually annotate the locations of visible objects in real scenes and provide the visible count and locations in our released datasets. Two examples are displayed in Fig. 12.
The results of this experiment can be found in Tab. 6. While the numbers for the 2D counting methods improve, it remains hard for them to distinguish similar objects clumped together. As a result, their performance re,ains much lower on these challenging scenes than on traditional 2D counting benchmarks.
Figure 12: Localization of visible objects. Left: $\mathcal{N}_{vis}=168$, Right: $\mathcal{N}_{vis}=96$
[…]
Visible objects only
BMNet+ [21] 0.51 0.65 51.28
SAM+CLIP [7,16] 0.57 0.81 50.84
All objects
BMNet+ [21] 0.93 0.98 131.44
SAM+CLIP [7,16] 0.94 0.99 124.31
Ours 0.36 0.06 53.31
Table 6: Counting visible and invisble objects separately.
@@ 11 Implementation details
We use the nerfstudio library [25] for 3D reconstruction, specifically the splatfacto method built on top of the gsplat library [32]. We thank the contributors of all the aforementioned libraries.